OpenStudy (anonymous):

solve the equation sqrt x+12=x

OpenStudy (angela210793):

is it all (x+12) under sqr?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{x+12}=x\]

OpenStudy (anonymous):

that is what it looks like

OpenStudy (angela210793):

raise to the 2nd power both sides x+12-x^2=0 do u know how to go further?

OpenStudy (anonymous):

somewhat

OpenStudy (angela210793):

D=b^2-4ac

OpenStudy (sasogeek):

do u own a graphing calculator?

OpenStudy (anonymous):

\[ x+12= x^{2}\]

OpenStudy (anonymous):

\[x^{2}-x-12=0\]

OpenStudy (anonymous):

\[\sqrt{x+12}\] \[\sqrt{x+12}^{2}=x ^{2} \] \[x+12=x^2\]

OpenStudy (anonymous):

(x-4)(x+3) x=4 and -3

OpenStudy (anonymous):

@mathballet how did you get that

OpenStudy (anonymous):

plug in to check for extraneous

OpenStudy (anonymous):

move the x^2 over to the other side correct

OpenStudy (anonymous):

brought everything to one side

OpenStudy (anonymous):

no dont move x^2

OpenStudy (anonymous):

\[-x^2 + x + 12 = 0;\ x \ge -12\]

OpenStudy (anonymous):

so move x+12

OpenStudy (anonymous):

You can move to either side.

OpenStudy (anonymous):

You get the same results.

OpenStudy (anonymous):

move x and 12

OpenStudy (anonymous):

Jenni, you're fine moving the x^2 if you'd rather. You get the same thing.

OpenStudy (anonymous):

its easier when factoring if u move x and 12

OpenStudy (anonymous):

Ah, yeah. if you're going to factor it's easier if you have a positive coefficient.

OpenStudy (anonymous):

ok so its (x+3)(x+4)?

OpenStudy (anonymous):

(x+3)(x-4) = 0

OpenStudy (anonymous):

(x-4)(x+3)

OpenStudy (anonymous):

but that would make -x-12+x^2=0

OpenStudy (anonymous):

if i move the x+12 to the other side

OpenStudy (anonymous):

xactly